Assume that a sample is used to estimate a population mean . Use the given confidence level and sample data to find the margin of error. Assume that the sample is a simple random sample and the population has a normal distribution. Round your answer to one more decimal place than the sample standard deviation. (please show work) 95% confidence; n = 51; X = 97; s = 202
Answer #1

Degrees of freedom = n-1

= 51 - 1

= 50

t critical value at 0.05 significance level for 50 df is 2.009

Margin of error = t * S / sqrt(n)

= 2.009 * 202 / sqrt(51)

= 56.8

Margin of error = 56.8 (Rounded one more decimal places than sample standard deviation)